OpenStudy (anonymous):

if a game ticket was 60.00 and than the price was increased by 2/5, what is the new cost of the ticket.?

OpenStudy (campbell_st):

well the increase is (60 * 2)/5 so the new price = 60 + increase

OpenStudy (imstuck):

In order to find the new price, you have to add 2/5 of 60 to 60. 2/5 = .4. .4 * 60 = $24, so 60 + 24 = $84.